OpenStudy (anonymous):

The time it takes a swimmer (in hours) to go a distance of 5 miles up and down a river is given by f(x)=4/x-1 hours and g(x)=3/x+1 hours respectively, where x is the speed of the swimmer (in miles per hour) in still water and (2≤ x ≤20). Find the expression for the total time it takes to go 5 miles up the river and come back to the same point. a 7x-1/x^2-1 b x+7/x^2-1 c 12/x^2-1 d 7x+1/x^2-1

OpenStudy (tkhunny):

Very hard to understand what you have written. Please evaluate: 8/4-1 = 9/(4-1) = Let's see if we are understanding each other.

OpenStudy (anonymous):

8/3 and 3

OpenStudy (tkhunny):

Okay, then we need to understand the Order of Operations a little better. 8/4-1 = (8/4) - 1 = 2 - 1 = 1 You must see that this is so.

OpenStudy (anonymous):

I think a is correct.

OpenStudy (anonymous):

no I think d

OpenStudy (tkhunny):

A is not correct as you have written it. Do you see that 8/4-1 = 1 and if you want 8/3 the parentheses are NOT optional? 8/(4-1) = 8/3 8/4-1 = 1

OpenStudy (anonymous):

I actually meant to write d.

OpenStudy (tkhunny):

D is also not correct as you have written it. You should have: f(x)=4/(x-1) hours and g(x)=3/(x+1) hours The parentheses are NOT optional. Distance = Rate * Time Trip Up the River: \(5 = x\cdot\dfrac{4}{x-1}\) Trip Down the River: \(5 = x\cdot\dfrac{3}{x+1}\)

OpenStudy (anonymous):

So c

OpenStudy (tkhunny):

C is not correct as you have written it. Why do you think so?